A 4.50 cm tall object is placed in front a convex mirror with a focal length of -6.25 cm. If the magnification is 1/4, what is the distance from the object to the mirror. Give your answer in centimeter (cm) and with 3 significant figures.
Answer #1

Here ,

height of object , ho = 4.50 cm

f = -6.25 m

magnification , m = 1/4

as m = -di/do

1/4 = - di/do

do = -4 * di ----(1)

Using lens formula

1/f = 1/di + 1/do

-1/6.25 = 1/di + 1/do ---(2)

solving 1 and 2

di = -4.7 cm

do = 18.75 cm

the object is placed at 18.75 cm from the lens